OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point shown and is parallel to the line shown.

OpenStudy (anonymous):

A. y = -4x - 6 B. y = 4x + 10 C. y = 1/4x + 5/2 D. y = -1/4x + 3/2

OpenStudy (anonymous):

I think it's B

OpenStudy (campbell_st):

you need to find the slope of the line it appears the line passes through the points (-4, -9) and (-3, -5) so the slope is (-9 - -5)/(-4 - -3) = 4 the point is (-2, 2) so the slope intercept form of the line is y = 4x + b substitute x = -2 and y = 2 to find b

OpenStudy (anonymous):

So the answer is A

OpenStudy (campbell_st):

nope... your answer is correct... and only answer B has a slope of 4....